Suppose a woman does 463 J of work and dissipates 8,854 J of heat in the process. (a) What is the decrease in her internal energy, assuming no change in temperature or consumption of food? (That is, there is no other energy transfer.) kcal (b) What is her efficiency? %
Answer #1

In the process, the total energy used = work + heat lost = 463 + 8854 = 9317 J

In kcal, energy used = 9317/(1000 x 4.18) = 2.23 kcal

As there is not change in temperature, there is no other change in internal energy.

Efficiency is the ratio of energy out / energy in = work / energy used = 463 / 9317 = 0.0497 = 4.97%.